OpenStudy (anonymous):

what about x^x = 27 is it possible to find out that x=3 by some method?

OpenStudy (anonymous):

Trial and error

OpenStudy (callisto):

xlogx = log27 xlogx = 3log3 x=3 ? huh.. it's tough!

OpenStudy (anonymous):

x^x=3^3 x=3

OpenStudy (callisto):

the only problem is that if you don't know 3^3 = 27, then what can be done?..... similarly, x^x = 287420498, what is x ?(surely that i know the answer but how can i find it if i don't know?)

OpenStudy (callisto):

ahhh,.... no no no log27 is not 3log9 and if you know xlogx=3log3 then you know the answer already...

OpenStudy (anonymous):

Maybe what you can do it factorize the number into its prime factors, that's what I can see

OpenStudy (anonymous):

xlogx=log27 xlogx=log(9*3) xlogx=2log3+log3 xlogx=3log3 xlogx-3log3=0 3xlog(x/3)=0 log(x/3)=0 logx-log3=0 logx=log3 x=3 but it should be avoided and concluded from xlogx=3log3

OpenStudy (callisto):

:) i see ~~ then can i just write xlogx=3log3 directly instead of ''xlogx=log(9*3) => xlogx=2log3+log3'' ?

OpenStudy (anonymous):

@Montreal: how did you get: xlogx-3log3=0 3xlog(x/3)=0

OpenStudy (anonymous):

youre very right.

OpenStudy (anonymous):

log x-log y= log(x/y) (if the logs have the same base)

OpenStudy (anonymous):

right, so howz xlogx - 3log3 = 3xlog(x/3)

OpenStudy (anonymous):

Its not right!

OpenStudy (anonymous):

so how do you group logs with factors

OpenStudy (anonymous):

xlogx-3log3 =log(x^x)-log(3^3) =log(x^x/3^3) This

OpenStudy (anonymous):

right thanks